MATH 110-2 Algebra Through History Practice with Different Bases September 6, 2019

Background

To practice using the ideas about different number systems introduced today, try the following problems/questions.

Questions

1. What base-10 number is represented by the base-8 number (231)_{8}?

2. What base-10 number is represented by the base-2 number (110011)_{2}?

3. How would you write (24)_{10} in base 8?

4. How would you write (24)_{10} in base 2? How is this related to your answer in question 3?